Identify the coefficients that would be used in the quadratic formula for this equation. 5x^2 = 4x
a. a = 5, b = 4, c = 1
b. a = -5, b = 4, c = 0
c. a = 5, b = 4, c = 0
d. a = 5, b = -4, c = 0

Answer by Theo(13342)

The standard form of the quadratic equation is:
ax^2 + bx + c = 0
-----
Your equation is 5x^2 = 4x
Substract 4x from both sides of this equation to get:
5x^2 - 4x = 0
Your a coefficient is 5
Your b coefficient is -4
Your c coefficient is 0
That would be selection d.
-----